# HENet: Hybrid Encoding for End-to-end Multi-task 3D Perception from Multi-view Cameras

### v2

@@ Abstract
Three-dimensional perception from multi-view cameras is a crucial component in autonomous driving systems, which involves multiple tasks like 3D object detection and bird’s-eye-view (BEV) semantic segmentation. To improve perception precision, large image encoders, high-resolution images, and long-term temporal inputs have been adopted in recent 3D perception models, bringing remarkable performance gains. However, these techniques are often incompatible in training and inference scenarios due to computational resource constraints. Besides, modern autonomous driving systems prefer to adopt an end-to-end framework for multi-task 3D perception, which can simplify the overall system architecture and reduce the implementation complexity. However, conflict between tasks often arises when optimizing multiple tasks jointly within an end-to-end 3D perception model. To alleviate these issues, we present an end-to-end framework named HENet for multi-task 3D perception in this paper. Specifically, we propose a hybrid image encoding network, using a large image encoder for short-term frames and a small image encoder for long-term temporal frames. Then, we introduce a temporal feature integration module based on the attention mechanism to fuse the features of different frames extracted by the two aforementioned hybrid image encoders. Finally, according to the characteristics of each perception task, we utilize BEV features of different grid sizes, independent BEV encoders, and task decoders for different tasks. Experimental results show that HENet achieves state-of-the-art end-to-end multi-task 3D perception results on the nuScenes benchmark, including 3D object detection and BEV semantic segmentation. The source code and models will be released at https://github.com/VDIGPKU/HENet.
@@ 1 Introduction
Efficient and accurate perception of the surrounding environment from multi-view cameras is essential for the autonomous driving system, which serves as the basis for subsequent trajectory prediction and motion planning tasks. A desirable 3D perception system should handle several tasks simultaneously, including 3D object detection and bird’s-eye-view (BEV) semantic segmentation. There has been an increasing emphasis on end-to-end multi-task frameworks, as such systems have the potential to streamline the overall architecture and alleviate implementation complexities.
@@ 4.3 Main Results
4.4 Analysis of Conflict in End-to-end Multi-task 3D Perception
4.5 Single Task Results
@@
3D Object Detection Results. We present the results of HENet for single 3D object detection task on the nuScenes val and test sets in Table 2 and 3, respectively. As shown in  Table 2, HENet surpasses all multi-view camera 3D object detection methods under different backbone and input resolution configurations, demonstrating the effectiveness of the proposed hybrid image encoding network and temporal feature integration module.
For online evaluation, Table 3 shows HENet achieves state-of-the-art 3D object detection results. To further demonstrate the effectiveness of HENet, we choose SparseBEV [26] as the baseline and incorporate it with HENet. We successfully implemented Hybrid Image Encoding of short-term ViT-L and long-term V2-99, proving the effectiveness of the Hybrid Image Encoding Network in reducing training costs. We achieve new state-of-the-art 3D detection results with 70.7 NDS and 64.5 mAP on nuScenes leaderboard (detection task/vision track).
@@
BEVFormerV2 [44] InternImage-XL 64.8 58.0 0.448 0.262 0.342 0.238 0.128
BEVDet-Gamma [12] †‡ Swin-B 66.4 58.6 0.375 0.243 0.377 0.174 0.123
SparseBEV [26] V2-99 67.5 60.3 0.425 0.239 0.311 0.172 0.116
StreamPETR-Large [37] ViT-L 67.6 62.0 0.470 0.241 0.258 0.236 0.134
Hop [53] ViT-L 68.5 62.4 0.367 0.249 0.353 0.171 0.131
Far3D [13] ViT-L 68.7 63.5 0.432 0.237 0.278 0.227 0.130
HENet_Sp (Ours) † ViT-L & V2-99 70.7 64.5 0.402 0.235 0.237 0.155 0.129
@@ 4.6 Ablation
Table 9: Multi-task results under different loss weights.
$\alpha_{depth}$ $\alpha_{cls}$ $\alpha_{bbox}$ $\alpha_{segveh}$ $\alpha_{segarea}$ $\alpha_{segdiv}$ NDS mIoU mIoUv mIoUa mIoUd
1 2.5 6 100 100 100 60.0 57.3 47.5 81.4 43.1
1 2.5 6 180 45 90 59.7 57.6 49.1 81.1 42.6
1 2.5 6 300 70 200 59.9 58.0 49.5 81.3 43.4
Loss Weights. Table 9 shows how we adjust the loss weight to obtain the results in Table 1. Specifically, we follow BEVDepth [18] to set $\alpha_{depth}=1$, $\alpha_{cls}=1.5$, and $\alpha_{box}=6$. We find that a larger segmentation loss weight achieves higher segmentation performance. Thus, we adjust the loss balance of each segmentation sub-task to obtain better overall segmentation results. The results show that carefully tuning loss weights can provide some improvement.

### v3

@@ Abstract
Three-dimensional perception from multi-view cameras is a crucial component in autonomous driving systems, which involves multiple tasks like 3D object detection and bird’s-eye-view (BEV) semantic segmentation. To improve perception precision, large image encoders, high-resolution images, and long-term temporal inputs have been adopted in recent 3D perception models, bringing remarkable performance gains. However, these techniques are often incompatible in training and inference scenarios due to computational resource constraints. Besides, modern autonomous driving systems prefer to adopt an end-to-end framework for multi-task 3D perception, which can simplify the overall system architecture and reduce the implementation complexity. However, conflict between tasks often arises when optimizing multiple tasks jointly within an end-to-end 3D perception model. To alleviate these issues, we present an end-to-end framework named HENet for multi-task 3D perception in this paper. Specifically, we propose a hybrid image encoding network, using a large image encoder for short-term frames and a small image encoder for long-term temporal frames. Then, we introduce a temporal feature integration module based on the attention mechanism to fuse the features of different frames extracted by the two aforementioned hybrid image encoders. Finally, according to the characteristics of each perception task, we utilize BEV features of different grid sizes, independent BEV encoders, and task decoders for different tasks. Experimental results show that HENet achieves state-of-the-art end-to-end multi-task 3D perception results on the nuScenes benchmark, including 3D object detection and BEV semantic segmentation.
@@ 1 Introduction
Efficient and accurate perception of the surrounding environment from multi-view cameras is essential for the autonomous driving system, which serves as the basis for subsequent trajectory prediction and motion planning tasks. A desirable 3D perception system should handle several tasks simultaneously, including 3D
@@
object detection and bird’s-eye-view (BEV) semantic segmentation. There has been an increasing emphasis on end-to-end multi-task frameworks, as such systems have the potential to streamline the overall architecture and alleviate implementation complexities.
@@ 4.1 Implementation Details
4.2 Dataset and Metrics
@@
4.3 Main Results
4.4 Analysis of Conflict in End-to-end Multi-task 3D Perception
@@
4.5 Single Task Results
3D Object Detection Results. We present the results of HENet for single 3D object detection task on the nuScenes val and test sets in Table 2 and 3, respectively. As shown in  Table 2, HENet surpasses all multi-view camera 3D object detection methods under different backbone and input resolution configurations, demonstrating the effectiveness of the proposed hybrid image encoding network and temporal feature integration module. For online evaluation, Table 3 shows HENet achieves state-of-the-art 3D object detection results.
@@ 5 Conclusion
Acknowledgments
This work was supported by National Key R&D Program of China (Grant No. 2022ZD0160305).